OpenStudy (anonymous):

can anyone check this answer? 7x^2+7y^2+56x+28y-35=0 x=(-4+sqrt21,0), (-4-sqrt21,0) y=(0,5),(0,1)

OpenStudy (anonymous):

can u explain a little more...tnx b4...

OpenStudy (anonymous):

This is the equation of a conic section. It will be a curve, it doesn't have a unique solution. What is your answer supposed to represent?

OpenStudy (anonymous):

Circle equation.. I just want to see if I got the intercepts correct

OpenStudy (anonymous):

It's not a circle.

OpenStudy (anonymous):

You can factor out the 7, multiply by 1/7 x^2 + y^2 + 8x + 4y - 35 = 0 the extra x and y terms make it not a circle.

OpenStudy (anonymous):

my problems says: A circle has and equation of: 7x^2+7y^2+56x+28y-35=0 find intercepts

OpenStudy (anonymous):

Yeah, sorry.

OpenStudy (anonymous):

@telliott99 no problem with ur work u just forgot to divide 35 by 7...:)

OpenStudy (anonymous):

and complete the squares. That's the key idea.

OpenStudy (anonymous):

equation of circle \( (x+4)^2+(y+2)^2=25 \)